# Evaluation of LLM Vulnerabilities to Being Misused for Personalized Disinformation Generation

### v1

@@ Appendix D Meta-evaluation Validation
For evaluation of model safety, agreement and disagreement with disinformation narratives, we have used the dataset [^6] https://github.com/kinit-sk/disinformation-capabilities of Vykopal et al. (2024) (we have used an average of the two human-annotations). For evaluation of text-generation model safety, we have used the answers for Q7 of the existing dataset, where the score of 1 represents refusal of the generation (i.e., activation of a safety-filter), the score of 2 represents disclaimer message accompanying the generated text, and the score of 3 is assigned otherwise. The score of 1 and 2 corresponds to “Yes” answer to our meta-evaluation questions regarding presence of safety-filter and disclaimer, respectively. Thus, a direct mapping is possible, resulting in the $\rho$ of 0.54. For evaluation of agreement and disagreement of the generated texts with the disinformation narrative, we have used answers for Q3 and Q4, respectively. Since they used a different scoring scale (score of 1 to 5), we have mapped “Yes” answer of meta-evaluation to the score of 5, “Partially” answer to the score of 3, and “No” answer to the score of 1, resulting in the $\rho$ of 0.69 for agreement and 0.53 for disagreement.
@@ Appendix F Supplementary Data
Table 6 summarizes results of the linguistic analysis with text-quality evaluation using various existing metrics described in Section 4.
Mean ($\pm$ Standard deviation)
Generator Characters Words Lines Sentences GRUEN LA OCQ
@@
Figure 12 illustrates the personalization-quality evaluation-scores distribution across the validation subset (balanced based on GPT-4o scores).
Table 7 contains detection performance evaluation using the existing dataset of Vykopal et al. (2024), where we provide AUC ROC (area under the curve of receiver operating characteristic) as a classification-threshold independent metric, classification Threshold calculated (calibrated) to maximize the difference between true positive and false positive rates (obtained from ROC curve), and MacroF1 representing macro average (due to class imbalance) of F1-score when the calculated thresholds are used for decisions. ROC curves of the detectors are illustrated in Figure 13. In Table 8, the detectability results of generated texts are provided per each generator. Differences between generators are in most cases statistically significant (paired t-test and Wilcoxon signed-rank test with p-values $\leq 0.05$) for Detection-Longformer and Binoculars detectors.
Detector AUC ROC Threshold MacroF1
Gemma-2-9-IT 0.97 0.9995 0.83
Detection-Longformer 0.97 0.9043 0.74
Binoculars 0.74 -0.9387 0.46
Table 7: Evaluation of the selected machine-generated detection methods.
Figure 13: Receiver operating characteristic curves of the selected detection methods.

### v2

@@ Appendix D Meta-evaluation Validation
For evaluation of model safety, agreement and disagreement with disinformation narratives, we have used the dataset [^6] https://github.com/kinit-sk/disinformation-capabilities of Vykopal et al. (2024) (we have used an average of the five human-annotations). For evaluation of text-generation model safety, we have used the answers for Q7 of the existing dataset, where the score of 1 represents refusal of the generation (i.e., activation of a safety-filter), the score of 2 represents disclaimer message accompanying the generated text, and the score of 3 is assigned otherwise. The score of 1 and 2 corresponds to “Yes” answer to our meta-evaluation questions regarding presence of safety-filter and disclaimer, respectively. Thus, a direct mapping is possible, resulting in the $\rho$ of 0.54. For evaluation of agreement and disagreement of the generated texts with the disinformation narrative, we have used answers for Q3 and Q4, respectively. Since they used a different scoring scale (score of 1 to 5), we have mapped “Yes” answer of meta-evaluation to the score of 5, “Partially” answer to the score of 3, and “No” answer to the score of 1, resulting in the $\rho$ of 0.69 for agreement and 0.53 for disagreement.
@@ Appendix F Supplementary Data
In Table 8, the detectability results of generated texts are provided per each generator. Differences between generators are in most cases statistically significant (paired t-test and Wilcoxon signed-rank test with p-values 0.05$\leq$ 0.05≤ 0.05) for Detection-Longformer and Binoculars detectors.